OpenStudy (anonymous):

Solve: x^4-34x^2=-225

OpenStudy (abdela25):

solve for "x" right ?

OpenStudy (anonymous):

x^4-34x^2 = -225. let y= x^2,then above equn becomes y^2-34y = - 225. y^2 - 34y + 225=0.

OpenStudy (anonymous):

then factor y^2 - 34y + 225=0

OpenStudy (anonymous):

it becomes (y-25)(y-9)

OpenStudy (anonymous):

now substitute the y by x^2 and solve for the x

OpenStudy (abdela25):

the answer would be x=+5 or x=+3 do you know how i got that or should me and @helpme1.2 guide you step by step if you want